OpenStudy (anonymous):

Use the Principle of Mathematical Induction to show that n^3+2n+6 is divisible by 3 for all natural numbers n.

OpenStudy (mathmate):

The proposition: n^3+2n+6 is divisible by 3 for all natural numbers n It's a three step process: 1. basis: for n=1, n^3+2n+6=1+2+6=9 =0 mod(3) 2. Assume the proposition is true for k, i.e. 3 | k^3+2k+6 3. Induction step: For n=k+1, we consider (k+1)^3+2(k+1)+6 =(k^3+3k^2+3k+1)+2k+2+6 =(k^3+3k+6) + 3k^2+1+2 =(k^3+3k+6) + 3(k^2+1) Now examine the expression. The first term has already been assumed to be true, so it's divisible by 3. The second term clearly has a factor of three, so it is also divisible by 3. We know that the sum of two quantities is divisible by 3 when each is divisible by three. Therefore the assumption is also true for n=k+1. Thus the proposition is true for all positive values of k.